If a electron is moving around a nucleus of atomic mass z find time period
 

Grade:12

1 Answers

Vikas TU
7 years ago
T= 2pi*r/nk
where k is the wavelength of the elctron associated.
Now also we know, r = 0.529*n^2/z
and k = h/mv
where m would be the mass of elctron and v = 0.02188 *z/n
putting all the values in T U would get the relaton as,
T = c*n^3/z^2
wer c would be constant U would get after solving.